Two stones are thrown with the same speed but making different angles with the horizontal. Their ranges are equal. If the angle of projection of one is ?/3 and its maximum height is y1 then the maximum height of the other will be


Grade:10

2 Answers

Kenny Toijam
10 years ago
If they have same horizontal ranges for different angles of projections,the angles are x/3 and (90°-x/3) The horizontal range for angle of projection x/3=u2 sin (2x/3) / g and range for 90°-(x/3)= u2 sin2[90°-(x/3)]/g=u2sin[180°-(2x/3)]/g But as the two ranges are equal as per the question,we get, x=(180*3)/4=135°.we know that y1=[u2*sin2(x/3)] / 2g for angle of projection (x/3) Let y2 be the maximum height for [90°-(x/3)] i.e y2= u2*sin2[ 90°-(x/3)] / 2g=y1*cot2(x/3)=y1*cot2(135/3)=y1.
keesom
10 years ago
ADITI ........................ for equal ranges the angle of projections are theta & 90- theta and after finding the angle u can easily find the maximum height .